OpenStudy (anonymous):

If r is the radius of a circle and d is its diameter, which of the following is an equivalent formula for the circumference C=2πr? A. C=2πd B. C=πD C. C=πr2 D. C=πdr please give me the best answer.

OpenStudy (anonymous):

B. C=πD

OpenStudy (anonymous):

thank you Mr.Professor ;)

OpenStudy (anonymous):

c=2*pi*r 2r=d so c=pi*d